OpenStudy (anonymous):

Write the expression as a single logarithm. log_340 – log_310

OpenStudy (zehanz):

There are rules you know! Rules for log that everyone with log probs should know: 1. log_a(x) + log_a(y) = log_a(xy) 2. log_a(x) - log_a(y) = log_a(x/y) 3. log_a(x^k) = k*log_a(x) 4. log_a(x)=log_b(x)/log_b(a) @Hero wrote your expression in the way you should have ;) Just apply rule nr...?

OpenStudy (anonymous):

Uh, loga4?

OpenStudy (anonymous):

I'm not stupid, I just don't have time to learn this stuff considering my work is due in about 8 hours.

OpenStudy (anonymous):

I know which rule it is. I read it when ZeHanz posted it. Is it Log_34?

OpenStudy (anonymous):

Theres a 3 n there not an a? and x/y is 40/10 which simplifies to 4

OpenStudy (zehanz):

I know you are not stupid and it was not my intention to imply that, sorry! Not having much time is the reason why you should know about the rules. You don't want to have to discover how it works everytime you get such a problem. The answer is:\[\log_{3} 4\]

OpenStudy (anonymous):

I did it wayyy up there ^ you just never read it :P

OpenStudy (anonymous):

Solve the exponential equation. 1/16=64^4x-3. What is the rule for that?

OpenStudy (anonymous):

(1/16)=64^(4x-3)

OpenStudy (zehanz):

I did read it, and just wanted to confirm you were right... 64 and 1/16 are both powers of 2, it could be helpful to write these numbers that way:\[\frac{ 1 }{ 16 }=\frac{ 1 }{ 2^4 }=2^{-4}\]\[64=2^6\]So:\[2^{-4}=(2^6)^{4x-3}=2^{24x-18}\]Now, because these are both equal powers of 2, the exponents should also be equal: -4=24x-18 Solve it to get x...

OpenStudy (zehanz):

Yes, even I understand it now!

OpenStudy (anonymous):

x=(14/24)?

OpenStudy (zehanz):

Well done, again!

OpenStudy (zehanz):

BTW, you may leave out the parentheses now, no need to overdo it :P

OpenStudy (anonymous):

Lol I'm parenthesis excited xD Okay so I know the rules for log problems but how do you solve log equations?

OpenStudy (anonymous):

I have to solve 1. 3 log 2x = 4 2. ln(5x + 7) = 8. 3. 5e^(2x=11)=30 And then I'll be done with this one.

OpenStudy (zehanz):

Always use:\[\log_{a}x=b \Leftrightarrow a^b=x \]This is the definition of log, in fact. With it , you can switch back and forth between logarithms and powers. 1. 3 * log(2x)=4, so log(2x)=4/3. Now remember log is log base 10, so according to the definition, log(2x)=4/3 can also be written as:\[10^{\frac{ 4 }{ 3 }}=x\] Now there is only some simplification of this last number to do (hint: it is (10^4)^(1/3)

OpenStudy (anonymous):

10^4 is 10000

OpenStudy (anonymous):

Can you solve it? Is it 21.54?

OpenStudy (anonymous):

@Hero can you check if 1 is 21.54?

OpenStudy (zehanz):

It is, but don't you have to give an exact answer?

OpenStudy (anonymous):

No. But the exact answer is 21.5443469. Can you help with the last two @ZeHanz :)

OpenStudy (zehanz):

I meant, \[10^{4/3}=\sqrt[3]{10^4}=\sqrt[3]{10^3 \cdot 10}=10\sqrt[3]{10}\]as an exact answer. I'll give the other two a try.

OpenStudy (anonymous):

I'm not in that advanced yet..

OpenStudy (zehanz):

Ok, sou you can round off. 2: \[\ln(5x+7)=8 \Leftrightarrow e^{8}=5x+7\] Solve for x and type into your calculator.

OpenStudy (zehanz):

Ok, sou you can round off. 2: \[\ln(5x+7)=8 \Leftrightarrow e^{8}=5x+7\] Solve for x and type into your calculator.

OpenStudy (anonymous):

How do I solve for x with those letters and such? I haven't studied any of this. :/

OpenStudy (zehanz):

Shouldn't "2x=11" read "2x-11"? Then:\[5e^{2x-11}=30\]So\[e^{2x-11}=6\]Uh oh, now the exponent is unknown... We can solve it (again) by looking at the rule \[\log_{a}x=b \Leftrightarrow a^b=x \]only we use it from right to left instead of the other way round:\[2x-11=\ln 6\]and away you go...

OpenStudy (anonymous):

My bad, it's 2x+11 **

OpenStudy (anonymous):

Is it 2x+11=ln6 then?

OpenStudy (zehanz):

Yes it is! Sorry, I was getting all excited about problem no. 3. If you can work with ordinary logs, such as \[\log_{3} x\]or \[\log_{10}x=\log x \]then you need to know that ln is the log with base 2.71828.... = e.

OpenStudy (anonymous):

So do you input what e is into the equation to solve?

OpenStudy (anonymous):

Or is the solution e^8=5x+7?

OpenStudy (anonymous):

Wait, x=(1/5) ln8 @ZeHanz ?

OpenStudy (zehanz):

Just use your calculator to get e^8, subtract 7 and divide by 5.